[figure: 9.7 cm]

HYPERFINE FIELD IN YIG WITH CHARGED SUBSTITUTION

J. Kohout^{a}, H. Štěpánková^{a}, J. Englich^{a}, P. Novák^{b}, M. Kučera^{a} and K. Nitsch^{b} aFaculty of Mathematics and Physics, Charles University, 180 00 Prague 8, V Holešovičkách 2, Czech Republic bInstitute of Physics, Academy of Sciences, Cukrovarnická 10, 16 253 Prague 6, Czech Republic

57Fe NMR spectra were measured on single crystal films of yttrium iron garnet with paired charged substitution (Ge^{4+}-Ca^{2+}) using spin echo technique at 4.2 K. A series of samples with various concentration of substitution was prepared by an epitaxial growth from BaO/B_{2}O_{3}/BaF_{2} on GGG substrate. Satellite lines corresponding to the iron ions having the Ge^{4+} and Ca^{2+} in their neighbourhood were identified. The sample with the germanium content of x(Ge)=0.03 and substrate orientation (110) was used for the further study: NMR spectra of this sample were measured in an external magnetic field of 0.1 T applied along several directions in the plane of the film. The modification of the isotropic part and the anisotropy tensor of the hyperfine interaction caused by the Ge^{4+} substituent were obtained from the experimental data. The results were compared with those previously reported for threevalent substitution (Ga^{3+}).

[figure: 13.4 cm]

Subject category :

2. Magnetic Films, Surfaces and Multilayers

Presentation mode : poster

Corresponding author : H. Štěpánková

Address for correspondence : Faculty of Mathematics and Physics Charles University 180 00 Prague 8 V Holešovičkách 2 Czech Republic

Email address : stepanko@mbox.troja.mff.cuni.cz